# 00footnotetext: Shared first author; Shared last author

### v3

@@ Introduction
Figure 1: Improving training sets leads to better models that are cheaper to train. Using DataComp-LM, we develop a high-quality dataset, DCLM-baseline, which we use to train models with state-of-the-art trade-off between compute and performance. We compare on both (left) a Core set of tasks and on (right) MMLU 5-shot. Specifically DCLM-baseline (orange) shows favorable performance relative to both close-source models (crosses) and other open-source datasets and models (circles). Models in this figure are from [161, 162, 4, 97, 22, 189, 154, 121, 43, 160, 68, 150, 130, 100, 10, 156].
@@
Beyond the lack of standardized benchmarks, another challenge for research on training data is that details about training sets are becoming increasingly rare, even for open weight models such as the Llama, Mistral, or Gemma models [161, 77, 154]. For all of these models, the training sets are not publicly available, and the corresponding model documentation only provides a coarse description of the respective training data, if any at all. As a result, it is currently unclear what ingredients constitute a state-of-the-art training set for language models.
To address these challenges, we introduce DataComp for Language Models (DCLM), the first benchmark for language model training data curation. In DCLM, researchers propose new training sets and data curation algorithms and then evaluate their datasets by training language models with a fixed training recipe on their data. By measuring the performance of the resulting model on downstream tasks, researchers can quantify the strengths and weaknesses of the corresponding training set.
@@
As a starting point for DCLM, we conduct 416 baseline experiments with different training sets and compute scales. Our experiments identify model-based filtering as a key component in an effective data curation pipeline. We also show that details of the filtering model can have a large impact on performance, ranging from 35% to 44% accuracy on MMLU 5-shot [71] at the 7B parameter scale (280B training tokens). Interestingly, a simple bigram classifier, combined with a carefully selected set of positive and negative examples, performs best among the classifiers we experimented with. In addition, we find that human quality judgments have only limited value in identifying high-quality training data.
@@
To collect large datasets for training LMs [30], researchers typically resort to web crawls, which can contain undesirable content that can be improved via curation. Most data curation efforts focus on methods for improving model performance [131, 30, 170, 128, 121, 150], including filtering by language [44, 131, 180, 86], heuristic-based filtering [59, 128, 121, 34, 150], quality filtering [170, 49, 99, 178, 139], data deduplication [3, 88] and mixing [177, 148, 6]. While prior work examines a limited set of filters, we conduct the largest public investigation of data curation, resulting in a strong DCLM-baseline dataset.
@@
As the scale of LMs has increased over the past years [161, 162, 4, 36, 73, 115, 128, 154], the community has curated larger datasets to match. Early works include the C4 dataset with 160 billion (B) tokens and The Pile [59] with 300B tokens. More recently, RefinedWeb [121] contains 600B tokens, Dolma [150] 3 trillion (T) tokens, FineWeb 15T tokens [122], and RedPajama-v2 30T tokens [43]. There are also large domain-specific datasets, such as the code-focused StackV2 with 900B tokens [101], as well as high-quality filtered subsets such as FineWeb-Edu [100] with 1.3T tokens. We include performance comparisons with various datasets in Figure 1 and examine FineWeb’s LightEval evaluation framework more closely in Appendix G. We release the largest pool of raw text data to date with 240T web-crawled tokens. We also release DCLM-baseline, a high-quality dataset from our pool that yields better models than prior datasets.
@@
Past work on benchmarking data improvements includes dataset distillation [46], curriculum learning [137], and transfer learning [5, 31]. In DataComp [57] and DataPerf [106], participants iterate on a dataset with a fixed model and training recipe for vision, vision-language, and speech tasks. The BabyLM challenge Loose track [167] focuses on efficient development of LMs with 125M to 220M parameters trained on 10M to 100M tokens. With a 200T token pool and 7B models, DCLM is the first large-scale data-centric benchmark for language models.
@@
This section describes the main components of DCLM. We start with DCLM-Pool, the raw text corpus underlying our benchmark (Section 3.1). We then develop the DCLM workflow, visualized in Figure 2: selecting a competition scale (Section 3.2), curating a dataset by filtering DCLM-Pool and potentially mixing in other sources (Section 3.3), training a model with fixed hyperparameters (Section 3.4), and evaluating the model to score the dataset (Section 3.5).
@@
Test set samples often contaminate language model training sets [48, 181, 51]; however, the effect of such samples on downstream performance remains largely unclear [88, 150, 115]. To allow researchers to better understand contamination, we release decontamination tooling instead of decontaminating DCLM-Pool directly. Our tools are based on Lee et al. [88] and allow participants to examine their datasets for overlap with our test sets. We ask all submissions to disclose a decontamination report and avoid using highly-contaminated data. For the highest scoring submissions, we plan to specifically evaluate them for contamination. In Section 4.6, we apply our tools to DCLM-Pool and evaluate whether contamination affects our models.
@@
To ensure DCLM is accessible to researchers with different compute constraints and to facilitate the study of scaling trends, we create different competition scales spanning three orders of compute magnitude (Table 1).Each scale (i.e., 400M-1x, 1B-1x, 1B-5x, 7B-1x, and 7B-2x) specifies the number of model parameters (e.g., 7B) and a Chinchilla multiplier (e.g., 1x). The number of training tokens for each scale is $20$$\times$ number of parameters $\times$ Chinchilla multiplier so that a multiplier of 1x corresponds to a compute allocation that Hoffmann et al. [73] found near-optimal.
A potential pitfall in our multi-scale design is that the ranking of data curation methods may change when increasing the compute scale. To better understand this concern, in Figure 3, we plot the performance of 10 methods at the 7B-1x scale as a function of their 400M-1x and 1B-1x performance. We find high rank correlation between the smaller 400M-1x, 1B-1x results and larger 7B-1x results (Pearson’s $r=0.885$ and $r=0.919$, respectively), suggesting better curation strategies at smaller scales transfer to larger scales. For more competition scale ablations, including experiments suggesting dataset improvements are largely orthogonal to training hyperparameters, see Appendix H.
@@
After choosing a scale, participants choose one of two tracks. (i) In the filtering track, participants propose algorithms to select training data from a candidate pool. We start with five pools, one for each scale in Table 1, which are random document subsets of DCLM-Pool. We restrict initial pool sizes by scale to encourage scalable filtering strategies and reflect realistic data download and storage constraints. (ii) In the mixing track, a submission combines documents from potentially many sources. For instance, participants can synthesize documents from DCLM-Pool, a custom crawl, Stack Overflow, and Wikipedia. Appendix C provides detailed rules for each track, and Appendix D describes our open-source, extensible tooling for executing filtering and mixing operations.
@@
Our full evaluation suite, based on LLM-Foundry [109], contains 53 downstream tasks suitable for base model evaluation (i.e., without fine-tuning): from question answering to open-ended generation formats, considering varied domains like coding, text-book knowledge, and common-sense reasoning. To evaluate data curation algorithms, we focus on three main performance metrics. First, we consider MMLU 5-shot accuracy [72], which is widely used to compare state-of-the-art models like GPT-4 [115] and Llama 3 70B [4]. Second, we propose Core centered accuracy, computed over a subset of 22 tasks (e.g., HellaSwag [186] and ARC-E [40]) that provide a low-variance signal even at small scales, linearly rescaling the accuracy per task so that 0 corresponds to random guessing and 1 corresponds to perfect accuracy. Finally, we report Extended centered accuracy, which averages the centered performance for all of our 53 tasks. For more metric details, see Appendix G.
@@
We now show how the DCLM workflow can lead to high-quality datasets and quantify the effect of data curation methods. This section describes the process of converting Common Crawl into our dataset, DCLM-baseline, as shown in Figure 4. We provide ablation experiments for each step along the way. We first evaluate open-source datasets as a starting point (Section 4.1). Next, we experiment with alternatives for several key phases of dataset construction: text extraction (Section 4.2), deduplication (Section 4.3), and model-based filtering (Section 4.4). We then experiment with mixing in high-quality sources (Section 4.5) and provide a contamination analysis (Section 4.6). In Section 5, we scale up this approach to train a 7B model for 2T tokens.
@@
Table 2: Comparison to existing datasets (7B-1x scale). Despite not mixing high-quality sources, RefinedWeb performs best.
@@
Text extraction is a common early processing step that pulls content from raw HTML. To understand the effect of this step, we compare three text extraction approaches: resiliparse, trafilatura (used by RefinedWeb), and the Common Crawl-provided WET files that contain pre-extracted text. We then apply RefinedWeb’s heuristic quality filters to each of the text extractions. In Table 3, we find both resiliparse and trafilatura improve Core by at least 2.5 points over the WET extraction. This is significant because most open source datasets, including C4, RedPajama, and Dolma-V1, use the WET extraction, which could partially explain their worse performance in Table 2. While resiliparse and trafilatura have similar downstream performance, resiliparse is 8×8$\times$8 × faster to run and hence more practical for large-scale processing. For more analysis, see Appendix J.
@@
Table 3: Comparison of text extractors (1B-1x scale). We apply three approaches for text extraction from HTML, process their output using the RefinedWeb heuristic quality filters, and evaluate the quality of models trained on the resulting datasets. We find stricter extractors such as resiliparse and trafilatura are superior to WET files provided by Common Crawl.
@@
Web-crawled datasets often contain many duplicate or near-duplicate data strings. Removing these duplicates from a training set serves the dual purpose of improving performance by reducing memorization [33, 88] and increasing data diversity. For deduplication, we explore MinHash [28], as part of a suffix array pipeline [88, 121], and near-duplicate Bloom filtering, which modifies an exact document and paragraph deduplication scheme [150]. We find that both approaches provide comparable downstream performance: within 0.2 Core percentage points at the 7B-2x scale. However, our modified Bloom filter approach scales more easily to datasets surpassing 10TB. We provide additional analysis in Appendix K.
@@
Recent literature [55, 150, 27] indicates that using learnable models as quality filters leads to downstream improvements. In this section, we investigate model-based filtering.
@@
We compare many strategies: 1) PageRank score filtering to retain documents based on how likely they are to be linked to other documents, 2) Semantic Deduplication (SemDedup) to remove documents with similar informational content [1], 3) linear classifiers fit on pre-trained BGE text embeddings [176], 4) AskLLM that prompts an LM to see if a document is helpful [139], 5) Perplexity filtering where we retain low perplexity sequences following CCNet [170], 6) Top-k average logits where we average the top-$k$ model logits over all words in a document to score how confident a model is that the correct words are within $k$ reasonable choices, and 7) fastText [81] binary classifiers to distinguish data quality. For training classifiers, we train on $\sim 400$k documents split equally between positive and negative classes. We experiment with different options for positive data and fix negative data as a sample from RefinedWeb. For the perplexity filtering and the top-k average logits strategies, we utilize a 154M parameter causal Transformer trained on a mix of English Wikipedia, the books subset of RedPajama v1, and peS2o  [149, 160]. We compare the aforementioned approaches in Table 4 and find that fastText-based filtering outperforms all other approaches. We next aim to understand how fastText training recipes affect its effectiveness as a data filtering network [55].
@@
To better understand the limits of fastText, we train several variants, exploring different choices for the reference data (i.e., the examples given positive labels), feature space, and filtering threshold, as shown in Table 5. For reference positive data, we considered commonly used sources like Wikipedia [59], OpenWebText2 [59], and RedPajama-books [160], following the reference data used for GPT-3 [30]. We also try a novel approach, using instruction-formatted data, drawing examples from OpenHermes 2.5 [157] (OH-2.5) and high-scoring posts from the r/ExplainLikeImFive (ELI5) subreddit. Overall, we find, when controlling for other hyperparameters, the fastText OH-2.5 +ELI5 approach gives a 3.5 percentage point lift on Core compared to the conventional choices. It is natural to ask whether using OH-2.5 data for filtering could preclude additional gains from instruction-tuning. In Appendix P, we show this is not the case, further suggesting the strength and compatibility of this approach with modern fine-tuning paradigms. Finally, we observe that using a fairly strict threshold, which keeps the top-10% of examples, helps over more permissive top-15% and top-20% thresholds. We further study the unintuitive behavior of dataset filtering and its connection to human judgment in Appendix M.
Table 5: fastText ablations (7B-1x scale). We ablate choices for the positive data (top) and threshold (bottom). ‘Dataset’ is positive set for fastText, while the negatives are randomly sampled from RefinedWeb. ‘Threshold’ is the percentile used for filtering based on fastText scores. “GPT-3 Approx” refers to a uniform mix of Wikipedia, OpenWebText2, and RPJ Books, as in [30].
@@
Researchers often combine Common Crawl (CC) with other data sources that are considered high-quality [160, 59, 65, 162] (e.g., Wikipedia, arXiv, Stack exchange, and peS2o [149]). Since DCLM participants can include additional data sources in the bring your own data track, we examined the potential benefits of adding high-quality sources to training sets derived from Common Crawl only. We compare a model trained on 100% filtered CC data to models trained with the mixing proportion from Llama 1 and RedPajama: 67% CC, and 33% from Wikipedia, Books, Stack exchange, arXiv, and Github. For the CC component, we consider different variants: a subset of our DCLM-baseline, RedPajama’s CC portion, RefinedWeb, and C4. The results in Table 6 show that mixing improves performance for the lower-performing CC subsets (C4, RedPajama-CC, and RefinedWeb). In the case of DCLM-baseline however, mixing actually hurts performance on average, which suggests it can be counterproductive given performant filtering. For additional mixing results, see Appendix L.
Table 6: Mixing high-quality sources with subsets of CommonCrawl (1B-1x scale). We evaluate the impact of mixing high-quality source (‘RPJ extras’) to various datasets derived from CommonCrawl, using the mixing ratios from Llama and RedPajama. Numbers in parentheses indicate the improvement/degradation due to mixing, compared to using only the base dataset.
@@
DCLM-baseline  [^1] The DCLM-baseline subset used here relied on an earlier version of our classifier and our pool. 31.1 29.5 ($-1.6$) 16.0 15.3 ($-0.7$)
@@
Here, we perform analysis to examine whether contamination of our pretraining data with our evaluation is an issue that influences our results. We focus on MMLU as our evaluation set of choice, given its popularity as a metric for language model performance at this scale.
As an experiment, we also attempt to detect and remove questions from MMLU that exist in DCLM-baseline. Our strategy is to flag training documents that contain the last sentence of a question from MMLU along with one of the corresponding options. For these flagged examples, we then remove all matched question and option strings. In order to improve recall, we opt to detect only the last sentence from each question, reducing the chance of missing questions due to formatting differences. Based on inspection, this also incurs many false positives. We then train a 7B-2x model with our DCLM-baseline without the detected MMLU overlap.
The results of this analysis can be seen in Table 7. We see that this removal of contaminated samples does not lead to a decrease in performance for our model. As such, we can see that our performance gains in MMLU are not caused by increased presence of MMLU in our dataset.
Table 7: MMLU overlap removal results. We remove overlaps detected with MMLU, in cases where a question and one of its options are detected in the text. We compare the performance between a model trained with and without this data removed, and see that there is no gain from increased contamination. This experiment is done at the 7B-2x scale.
Dataset MMLU
DCLM-baseline 51.8
DCLM-baseline (MMLU removed) 52.7
We also apply the above removal strategy on Dolma-V1.7 [150] and FineWeb-Edu [100], to measure what the contamination differences are between DCLM-baseline and those datasets. The results can be seen in Table 8. We see that our DCLM-baseline has roughly similar contamination stats as other high performing datasets according to this analysis. We provide further contamination analysis that extends to our entire benchmark suite in Appendix N.
Dataset Percentage of samples flagged
DCLM-baseline 0.007%
Dolma-V1.7 0.001%
FineWeb-Edu 0.009%
@@
Here, we test if datasets that perform well on the DCLM benchmark also maintain their strength with an order of magnitude more compute. To ensure our trained model is broadly useful, including for math and coding tasks, we combine our 3.8T DCLM-baseline with the StarCoder [90] and ProofPile2 [14] data to arrive at a 4.1T token dataset. We train a 7B model for 2.5T tokens on this dataset with the same hyperparameters as our largest competition scale except for two separate cool-downs phase for the 200B and 270B tokens on a modified distribution that was 70% DCLM-baseline with a tighter fastText threshold, and 30% math datasets (see Appendix P). We then take a “model soup" of these two separate cool-downs[173]. We then adopt the continual pre-training methodology from Pouransari et al. [126] for 100B tokens on the same distribution to increase the context length from 2048 to 8192, we provide more details on this procedure in Section P.2.
In Table 9, we show that our model outperforms all 7B models trained on public training sets and approaches closed-data models trained for more tokens such as Llama-8B, Mistral-7B, and Gemma-7B. Additionally, in Appendix O, we show that our model achieves strong instruction-tuning performance. After instruction tuning on publicly available IT datasets, our model maintains most of its benchmark performance and achieves an AlpacaEval2.0 LC Win-rate of 16.6, which outperforms Gemma-Instruct (10.4), while approaching the strong performance of Mistral-v0.2-7B (17.1) and Llama3-Instruct (22.9).
Table 9: State-of-the-art comparison (beyond 7B-2x scale). We compare our final model with others in the 7–8B parameter regime. Our DCLM-baseline dataset yields a model that outperforms models trained on open datasets and is competitive with models trained on private datasets.

### v4

# DataComp-LM: In search of the next generation of training sets for language models

@@ Introduction
Figure 1: Improving training sets leads to better models that are cheaper to train. Using DataComp-LM, we develop a high-quality dataset, DCLM-baseline, which we use to train models with state-of-the-art trade-offs between compute and performance. We compare on both (left) a Core set of tasks and on (right) MMLU 5-shot. Specifically DCLM-baseline (orange) shows favorable performance relative to both closed-source models (crosses) and other open-source datasets and models (circles). Models in this figure are from [169, 170, 4, 103, 22, 198, 162, 128, 46, 168, 73, 157, 137, 106, 10, 164, 103]. Table 33 provides a table version of this figure.
@@
Beyond the lack of standardized benchmarks, another challenge for research on training data is that details about training sets are becoming increasingly rare, even for open weight models such as the Llama, Mistral, or Gemma models [169, 83, 162]. For all of these models, the training sets are not publicly available, and the corresponding model documentation only provides a coarse description of the respective training data, if any at all. As a result, it is currently unclear what ingredients constitute a state-of-the-art training set for langauge models.
To address these challenges, we introduce DataComp for Language Models (DCLM), the first large-scale benchmark for language model training data curation. In DCLM, researchers propose new training sets and data curation algorithms and then evaluate their datasets by training LMs with a fixed training recipe on their data. By measuring the performance of the resulting model on downstream tasks, researchers can quantify the strengths and weaknesses of the corresponding training set.
@@
As a starting point for DCLM, we conduct 416 baseline experiments with different training sets and compute scales. Our experiments identify model-based filtering as a key component for effective data curation. We also show that details of the filtering model can have a large impact on performance, ranging from 35% to 44% accuracy on MMLU 5-shot [77] at the 7B parameter scale (280B training tokens). Interestingly, a simple bigram classifier, combined with a carefully selected set of positive and negative examples, performs best among the classifiers we experimented with. In addition, we find that human quality judgments have only limited value in identifying high-quality training data.
@@
To collect large datasets for training LMs [31], researchers typically resort to web crawls. However, these crawls often contain considerable amounts of undesirable content, necessitating curation to develop high-quality training data. Most data curation efforts focus on methods for improving model performance [138, 31, 178, 135, 128, 157], including filtering by language [47, 138, 189, 92], heuristic-based filtering [63, 135, 128, 37, 157], quality filtering [178, 53, 105, 187, 146], data deduplication [3, 94] and mixing [186, 155, 6]. While prior work examines a limited set of filters, we conduct the largest public investigation of data curation, resulting in a strong DCLM-baseline dataset. Since the initial release of DCLM, newer works such as WebOrganizer [180], Nemotron-CC [159], and Olmo-2 [121] have also built upon our benchmark or curation strategies to further advance the state-of-the-art for LLM pre-training datasets.
@@
As the scale of LMs has increased over the past years [169, 170, 4, 39, 79, 122, 135, 162], the community has curated larger datasets to match. Early works include the C4 dataset with 160 billion (B) tokens and The Pile [63] with 300B tokens. More recently, RefinedWeb [128] contains 600B tokens, Dolma [157] 3 trillion (T) tokens, FineWeb 15T tokens [129], and RedPajama-v2 30T tokens [46]. There are also large domain-specific datasets, such as the code-focused StackV2 with 900B tokens [107], as well as high-quality filtered subsets such as FineWeb-Edu [106] with 1.3T tokens. We include performance comparisons with various datasets in Figure 1 and examine FineWeb’s LightEval evaluation framework more closely in Appendix G. We release the largest pool of raw text data to date with 240T web-crawled tokens. We also release DCLM-baseline, a 3.8T token high-quality dataset from our pool that yields better models than prior datasets.
@@
Past work on benchmarking data improvements includes dataset distillation [50], curriculum learning [144], and transfer learning [5, 32]. In DataComp [61] and DataPerf [112], participants iterate on a dataset with a fixed model and training recipe for vision, vision-language, and speech tasks. For LMs, the Data-Juicer [36] effort includes benchmarks for cleaning and mixing fine-tuning data while the BabyLM challenge Loose track [175] focuses on efficient development of 125M to 220M parameter LMs pretrained on 10M to 100M tokens. With a 240T token pool and 7B models, DCLM is the largest data-centric benchmark for language models.
@@
This section describes the main components of DCLM. We start with DCLM-Pool, the raw text corpus underlying our benchmark (Section 3.1). We then describe the DCLM workflow, visualized in Figure 2: selecting a competition scale (Section 3.2), curating a dataset by filtering DCLM-Pool and potentially mixing in other sources (Section 3.3), training a model with fixed hyperparameters (Section 3.4), and evaluating the model to score the dataset (Section 3.5).
@@
Test set samples often contaminate language model training sets [52, 190, 55]; however, the effect of such samples on downstream performance remains largely unclear [94, 157, 122]. To allow researchers to better understand contamination, we release decontamination tooling instead of decontaminating DCLM-Pool directly. First, as used in Section 4.6, we implement our own decontamination process for two popular tasks, MMLU and Hellaswag [78, 195]. Second, we provide tooling based on Lee et al. [94] to examine datasets for overlap with all of our test sets. We ask all submissions to disclose a decontamination report and avoid using highly-contaminated data. For the highest scoring submissions, we plan to specifically evaluate them for contamination.
@@
To ensure DCLM is accessible to researchers with different compute constraints and to facilitate the study of scaling trends, we create different competition scales spanning three orders of compute magnitude (Table 1). Each scale (i.e., 400M-1x, 1B-1x, 3B-1x, 7B-1x, and 7B-2x) specifies the number of model parameters (e.g., 7B) and a Chinchilla multiplier (e.g., 1x). The number of training tokens for each scale is $20$$\times$ number of parameters $\times$ Chinchilla multiplier so that a multiplier of 1x corresponds to a compute allocation that Hoffmann et al. [79] found near-optimal.
A potential pitfall in our multi-scale design is that the ranking of data curation methods may change when increasing the compute scale. To better understand this concern, in Figure 3, we plot the performance of 10 methods at the 7B-1x scale as a function of their performance at smaller scales. We find high rank correlation between the results for smaller scales (400M-1x, 1B-1x, 3B-1x) and those for the larger 7B-1x scale (Pearson’s $r=0.838$, $r=0.956$, $r=0.982$ respectively), suggesting better curation strategies at smaller scales transfer to larger scales. For more competition scale ablations, including experiments suggesting dataset improvements are largely orthogonal to training hyperparameters, see Appendix H.
@@
After choosing a scale, participants choose one of two tracks. (i) In the filtering track, participants propose algorithms to select training data from a candidate pool. We start with five pools, one for each scale in Table 1, which are random document subsets of DCLM-Pool. We restrict initial pool sizes by scale to encourage scalable filtering strategies and reflect realistic data download and storage constraints. (ii) In the mixing track, a submission may combine documents from DCLM-Pool with those from any external sources (i.e., outside of Common Crawl). For instance, participants can add in documents directly curated from StackExchange and Wikipedia data dumps (as is done by RPJ [168]) or perhaps even their own custom crawl. Appendix C provides more detailed rules for each track, and Appendix D describes our extensible open-source tooling for executing filtering and mixing operations.
@@
Our full evaluation suite, based on LLM-Foundry [115], contains 53 downstream tasks suitable for base model evaluation (i.e., without finetuning): from question answering to open-ended generation formats, considering varied domains like mathematics, text-book knowledge, and common-sense reasoning. To evaluate data curation algorithms, we focus on three main performance metrics. First, we consider MMLU 5-shot accuracy [78], which is widely used to compare state-of-the-art models like GPT-4 [122] and Llama 3 70B [4]. Second, we propose the Core centered accuracy, computed over a subset of 22 tasks (e.g., HellaSwag [195] and ARC-E [43]) that provide a low-variance signal even at small scales, linearly rescaling the accuracy per task so that 0 corresponds to random guessing and 1 corresponds to perfect accuracy. Finally, we report the Extended centered accuracy, which averages the centered performance for all of our 53 tasks. For more metric details, see Appendix G.
@@
We now show how the DCLM workflow can lead to high-quality datasets and quantify the effects of data curation methods. This section describes the process of converting Common Crawl into our dataset, DCLM-baseline, as shown in Figure 4. We provide ablation experiments for each step along the way. We first evaluate open-source datasets as a starting point (Section 4.1). Next, we experiment with alternatives for several key phases of dataset construction: text extraction (Section 4.2), deduplication (Section 4.3), and model-based filtering (Section 4.4). We then experiment with mixing in high-quality sources (Section 4.5) and provide a contamination analysis (Section 4.6). In Section 5, we scale up this approach to train a 7B model for 2T tokens.
@@
Table 2: Comparison to existing datasets (7B-1x scale). Despite not mixing high-quality sources (unlike Dolma-V1 and RedPajama), RefinedWeb performs best.
@@
Text extraction is a crucial early processing step that pulls content from raw HTML. To understand the effect of this step, we compare three extraction approaches: resiliparse, trafilatura (used by RefinedWeb), and the Common Crawl-provided WET files that contain pre-extracted text. We then apply RefinedWeb’s heuristic quality filters to each of the text extractions. In Table 3, we find both resiliparse and trafilatura improve Core by at least 2.5 points over the WET extraction. This is significant because most open source datasets, including C4, RedPajama, and Dolma-V1, use WET files, which could partially explain their worse performance in Table 2. While resiliparse and trafilatura have similar downstream performance, resiliparse is 8×8$\times$8 × faster to run and hence more practical for large-scale processing. For more analysis, see Appendix K.
@@
Table 3: Comparison of text extractors (1B-1x scale). We apply three approaches for text extraction from HTML, process their output using the RefinedWeb heuristic filters, and evaluate the models trained on the resulting datasets. We find stricter extractors such as resiliparse and trafilatura are superior to WET files provided by Common Crawl.
@@
Web-crawled datasets often contain many duplicate or near-duplicate data strings. Removing these duplicates serves the dual purpose of improving performance by reducing memorization [35, 94] and increasing data diversity. For deduplication, we explore MinHash [29], as part of a suffix array pipeline [94, 128], and near-duplicate Bloom filtering, which modifies an exact document and paragraph deduplication scheme [157]. We find that both approaches provide comparable downstream performance: within 0.2 Core percentage points at the 7B-2x scale. However, our modified Bloom filter scales more easily to datasets surpassing 10TB. We provide additional analysis in Appendix L.
@@
Recent literature [59, 157, 28] indicates that training models to serve as quality filters leads to downstream improvements. In this section, we investigate different model-based filtering approaches.
@@
We compare many strategies: 1) PageRank score filtering to retain documents based on how likely they are to be linked to other documents, 2) Semantic Deduplication (SemDedup) to remove documents with similar informational content [1], 3) linear classifiers fit on pre-trained BGE text embeddings [185], 4) AskLLM which prompts an LM to see if a document is helpful [146], 5) Perplexity filtering where we retain low perplexity sequences following CCNet [178], 6) Top-k average logits where we average the top-$k$ model logits over all words in a document to score how confident a model is that the correct words are within $k$ reasonable choices, and 7) fastText [87] binary classifiers to distinguish data quality. For training classifiers, we train on $\sim 400$k documents split equally between positive and negative classes. We experiment with different options for positive data and fix negative data as a random sample from a version of our RefinedWeb reproduction. For the perplexity filtering and the top-k average logits strategies, we utilize a 154M parameter causal Transformer trained on a mix of English Wikipedia, the books subset of RedPajama-v1, and peS2o  [156, 168] (see Appendix J for more implementation details). We compare the aforementioned approaches in Table 4 and find that fastText-based filtering outperforms all other approaches. We next aim to understand how different fastText training recipes affect its effectiveness as a data filtering network [59].
@@
To better understand the limits of fastText, we train several variants, exploring different choices for the reference data (i.e., the examples given positive labels), filtering threshold, and feature space as shown in Tables 5 and 14. For reference data, we considered commonly used sources like Wikipedia [63], OpenWebText2 [63], and RedPajama-books [168], following the reference data used for GPT-3 [31]. We also try a novel approach, using instruction-formatted data, drawing examples from OpenHermes 2.5 [165] (OH-2.5) and high-scoring posts from the r/ExplainLikeImFive (ELI5) subreddit. Overall, we find, when controlling for other hyperparameters, the fastText OH-2.5 +ELI5 approach gives a 3.5 percentage point lift on Core compared to the other more conventional choices. It is natural to ask whether using OH-2.5 data for filtering could preclude additional gains from instruction-tuning. In Appendix Q, we show this is not the case, further suggesting the strength and compatibility of this approach with modern finetuning paradigms. Finally, we observe that using a fairly strict threshold, which keeps the top-10% of examples, helps over more permissive top-15% and top-20% thresholds. We further study the unintuitive behavior of dataset filtering and its connection to human judgment in Appendix N.
Table 5: fastText ablations (7B-1x scale). We ablate choices for the positive data (top) and threshold (bottom). ‘Dataset’ is the positive set, while the negatives are randomly sampled our RefinedWeb reproduction. ‘Threshold’ is the percentile used for filtering based on fastText scores. “GPT-3 Approx” refers to a mix of Wikipedia, OpenWebText2, and RPJ Books, as in [31].
@@
Often, Common Crawl (CC) is combined with other data sources that are considered high-quality [168, 63, 70, 170] (e.g., Wikipedia, StackExchange, and peS2o [156]). Since DCLM participants can include additional data sources in our mixing track, we examined the potential benefits of adding high-quality sources to training sets derived from Common Crawl only. We compare a model trained on 100% filtered CC data to models trained with the mixing ratios from Llama 1 and RedPajama: 67% CC, and 33% from Wikipedia, Books, Stack exchange, arXiv, and Github. For the CC component, we consider different variants: a subset of our DCLM-baseline, RedPajama’s CC portion, RefinedWeb, and C4. The results in Table 6 show that mixing improves performance for the lower-performing CC subsets (C4, RedPajama-CC, and RefinedWeb). In the case of DCLM-baseline however, mixing actually hurts performance on average, which suggests it can be counterproductive given performant filtering. For additional mixing results, see Appendix M.
Table 6: Mixing high-quality sources with subsets of CommonCrawl (1B-1x scale). We evaluate the impact of mixing high-quality sources (‘RPJ extras’) with various datasets derived from CommonCrawl, using the mixing ratios from Llama/RPJ. Numbers in parentheses indicate the gain or loss in performance due to mixing compared to using only the base dataset.
@@
DCLM-baseline 31.1 29.9 ($-1.2$) 16.0 15.0 ($-1.0$)
@@
Here, we examine whether contamination of our pretraining data with evaluation data influences our results for DCLM-baseline. We focus on MMLU and Hellaswag as our evaluation sets of choice, given their popularity as metrics for language model performance at the 7B scale.
Specifically, we attempt to remove examples from these two test sets that exist in DCLM-baseline. For both, we flag pages that contain the question text along with at least one of the corresponding answer options for a given example. For these flagged pages, we then remove all matched question and option strings. In order to improve recall for MMLU, which contains some long passage-based questions, we opt to detect only the last sentence from each question, reducing the chance of missing questions due to formatting differences. Based on inspection, this still incurs many false positives. We then train a 7B-2x model with our DCLM-baseline without the detected overlaps. As seen in Table 7, this does not lead to decreases in model performance, so our performance gains on these two tasks are not likely to be caused by increased presence of their test examples in our dataset.
Table 7: MMLU and Hellaswag overlap removal (7B-2x scale). We remove overlaps detected with MMLU and Hellaswag, in cases where a question and one of its options are detected. We compare models trained before and after this decontamination step, and see that performance does not fall.
Dataset $E$ = MMLU $E$ = Hellaswag
DCLM-baseline 51.8 77.9
DCLM-baseline ($E$ removed) 52.7 78.4
We also apply the above removal strategy for MMLU on Dolma-V1.7 [157] and FineWeb-Edu [106]. These results can be seen in Table 25 in Appendix O, from which we observe that DCLM-baseline has roughly similar contamination stats as these other high performing datasets. We also provide further analysis that extends to our entire evaluation suite in Appendix O.
@@
Here, we test if datasets that perform well on the DCLM benchmark also maintain their strength with an order of magnitude more compute. To ensure our trained model is broadly useful, including for math and coding tasks, we combine our 3.8T DCLM-baseline with the StarCoder [96] and ProofPile2 [14] datasets to arrive at a 4.1T token dataset. We train a 7B model for 2.5T tokens on this dataset with the same hyperparameters as our largest competition scale except for two separate cool-downs phase for the 200B and 270B tokens on a modified distribution that was 70% DCLM-baseline with a tighter fastText threshold, and 30% math datasets (see Appendix Q). We then take a “model soup" of these two separate cool-downs[182]. Finally, we adopt the continual pretraining methodology from Pouransari et al. [133] for 100B tokens on the same distribution to increase the context length from 2048 to 8192 (see Section Q.2).
In Table 8, we show that our model outperforms all 7B models trained on public training sets and approaches closed-data models trained for more tokens such as Llama-8B, Mistral-7B, and Gemma-7B. Additionally, in Appendix P, we show that our model achieves strong instruction-tuning (IT) performance. After instruction tuning on publicly available IT datasets, our model maintains most of its benchmark performance and achieves an AlpacaEval2.0 LC Win-rate of 16.6, which outperforms Gemma-Instruct (10.4), while approaching the strong performance of Mistral-v0.2-7B (17.1) and Llama3-Instruct (22.9). Finally, in Section Q.3, we show results from training a 1B model on 4.3T tokens from DCLM-baseline, StarCoder and ProofPile2 combined, resulting in a strong, small model that outperforms prior small models including Gemma-2B and Qwen2-1.5B.
Table 8: State-of-the-art comparison (beyond 7B-2x scale). We compare our final model with other 7–8B parameter models. DCLM-baseline yields a model that outperforms models trained on open datasets and is competitive with models trained on private datasets.